# Simple is Effective: The Roles of Graphs and Large Language Models in Knowledge-Graph-Based Retrieval-Augmented Generation

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate strong reasoning abilities but face limitations such as hallucinations and outdated knowledge. Knowledge Graph (KG)-based Retrieval-Augmented Generation (RAG) addresses these issues by grounding LLM outputs in structured external knowledge from KGs. However, current KG-based RAG frameworks still struggle to optimize the trade-off between retrieval effectiveness and efficiency in identifying a suitable amount of relevant graph information for the LLM to digest. We introduce SubgraphRAG, extending the KG-based RAG framework that retrieves subgraphs and leverages LLMs for reasoning and answer prediction. Our approach innovatively integrates a lightweight multilayer perceptron with a parallel triple-scoring mechanism for efficient and flexible subgraph retrieval while encoding directional structural distances to enhance retrieval effectiveness. The size of retrieved subgraphs can be flexibly adjusted to match the query’s need and the downstream LLM’s capabilities. This design strikes a balance between model complexity and reasoning power, enabling scalable and generalizable retrieval processes. Notably, based on our retrieved subgraphs, smaller LLMs like Llama3.1-8B-Instruct deliver competitive results with explainable reasoning, while larger models like GPT-4o achieve state-of-the-art accuracy compared with previous baselines—all without fine-tuning. Extensive evaluations on the WebQSP and CWQ benchmarks highlight SubgraphRAG’s strengths in efficiency, accuracy, and reliability by reducing hallucinations and improving response grounding.
@@ 3.2 Prompting-Based LLM Reasoning
Regarding the size $K$ of the retrieved subgraph, while increasing $K$ in principle improves the coverage of relevant information, it also incurs higher costs/latency for LLM reasoning and risks introducing more irrelevant information that may ultimately hurt LLM reasoning (Xu et al., 2024; Liu et al., 2024b). Different LLMs are inherently equipped with different sized context window and also exhibit distinct capabilities in reasoning over long-context retrieval results (Dubey et al., 2024). As such, although the training of SubgraphRAG retriever is LLM-agnostic, the size $K$ needs to be properly selected per LLM and cost/latency constraint. In Section 4.2, we empirically verify that more powerful LLMs can benefit from incorporating a larger-sized retrieved subgraph, demonstrating the benefit of size-adjustable subgraph retrieval in SubgraphRAG.
@@ 4.2 KGQA Results (Q3 & Q4)
Dataset Method Samples w/ Ans Entities in KG Samples w/o Ans Entities in KG No Ans SamplesTotal $\frac{\text{No Ans Samples}}{\text{Total Samples}}$ Correct AnsTotal $\frac{\text{Correct Ans}}{\text{Total Ans}}$ Wrong AnsTotal $\frac{\text{Wrong Ans}}{\text{Total Ans}}$ Correct Ans (NR) Correct $\frac{\text{Correct Ans (NR) }}{\text{Correct Ans}}$ Wrong Ans (NR)Wrong $\frac{\text{Wrong Ans (NR)}}{\text{Wrong Ans}}$ No Ans Samples Total $\frac{\text{No Ans Samples }}{\text{Total Samples}}$ Correct Ans (NR)Total $\frac{\text{Correct Ans (NR)}}{\text{Total Ans}}$ Wrong Ans (NR)Total $\frac{\text{Wrong Ans (NR)}}{\text{Total Ans}}$ Wrong Ans (R)Total $\frac{\text{Wrong Ans (R)}}{\text{Total Ans}}$ WebQSP RoG-Joint 0/1559 = 0% 5601/10206 = 55% 4605/10206 = 45% 170/5601 = 3% 1178/4605 = 26% 0/69 = 0% 32/126 = 25% 24/126 = 19% 70/126 = 56% RoG-Sep 0/1559 = 0% 5534/12401 = 45% 6867/12401 = 55% 327/5534 = 6% 2494/6867 = 36% 0/69 = 0% 48/162 = 30% 71/162 = 44% 43/162 = 27% SubgraphRAG (Llama3.1-8B) 29/1559 = 2% 4453/5850 = 76% 1397/5850 = 24% 59/4453 = 1% 84/1397 = 6% 13/69 = 19% 20/107 = 19% 16/107 = 15% 71/107 = 66% SubgraphRAG (GPT-4o-mini) 12/1559 = 1% 6209/8011 = 78% 1802/8011 = 22% 37/6209 = 1% 126/1802 = 7% 7/69 = 10% 35/83 = 42% 17/83 = 20% 31/83 = 37% CWQ RoG-Joint 0/2848 = 0% 3058/6709 = 46% 3651/6709 = 54% 305/3058 = 10% 1293/3651 = 35% 0/683 = 0% 531/2729 = 19% 1835/2729 = 67% 363/2729 = 13% RoG-Sep 0/2848 = 0% 2946/6129 = 48% 3183/6129 = 52% 341/2946 = 12% 1295/3183 = 41% 0/683 = 0% 512/2620 = 20% 1840/2620 = 70% 268/2620 = 10% SubgraphRAG (Llama3.1-8B) 210/2848 = 7% 2611/5028 = 52% 2417/5028 = 48% 21/2611 = 1% 98/2417 = 4% 199/683 = 29% 119/1052 = 11% 114/1052 = 11% 819/1052 = 78% SubgraphRAG (GPT-4o-mini) 203/2848 = 7% 3011/5205 = 58% 2194/5205 = 42% 30/3011 = 1% 159/2194 = 7% 137/683 = 20% 219/953 = 23% 170/953 = 18% 564/953 = 59%

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate strong reasoning abilities but face limitations such as hallucinations and outdated knowledge. Knowledge Graph (KG)-based Retrieval-Augmented Generation (RAG) addresses these issues by grounding LLM outputs in structured external knowledge from KGs. However, current KG-based RAG frameworks still struggle to optimize the trade-off between retrieval effectiveness and efficiency in identifying a suitable amount of relevant graph information for the LLM to digest. We introduce SubgraphRAG, extending the KG-based RAG framework that retrieves subgraphs and leverages LLMs for reasoning and answer prediction. Our approach innovatively integrates a lightweight multilayer perceptron with a parallel triple-scoring mechanism for efficient and flexible subgraph retrieval while encoding directional structural distances to enhance retrieval effectiveness. The size of retrieved subgraphs can be flexibly adjusted to match the query’s need and the downstream LLM’s capabilities. This design strikes a balance between model complexity and reasoning power, enabling scalable and generalizable retrieval processes. Notably, based on our retrieved subgraphs, smaller LLMs like Llama3.1-8B-Instruct deliver competitive results with explainable reasoning, while larger models like GPT-4o achieve state-of-the-art accuracy compared with previous baselines—all without fine-tuning. Extensive evaluations on the WebQSP and CWQ benchmarks highlight SubgraphRAG’s strengths in efficiency, accuracy, and reliability by reducing hallucinations and improving response grounding. Our implementation is available at https://github.com/Graph-COM/SubgraphRAG.
@@ 3.2 Prompting-Based LLM Reasoning
Regarding the size $K$ of the retrieved subgraph, while increasing $K$ in principle improves the coverage of relevant information, it also incurs higher costs/latency for LLM reasoning and risks introducing more irrelevant information that may ultimately hurt LLM reasoning (Xu et al., 2024; Liu et al., 2024b). Different LLMs are inherently equipped with different sized context window and also exhibit distinct capabilities in reasoning over long-context retrieval results (Dubey et al., 2024; Leng et al., 2024). As such, although the training of SubgraphRAG retriever is LLM-agnostic, the size $K$ needs to be properly selected per LLM and cost/latency constraint. In Section 4.2, we empirically verify that more powerful LLMs can benefit from incorporating a larger-sized retrieved subgraph, demonstrating the benefit of size-adjustable subgraph retrieval in SubgraphRAG.
@@ 4.2 KGQA Results (Q3 & Q4)
Dataset Method Samples w/ Ans Entities in KG Samples w/o Ans Entities in KG No Ans SamplesTotal $\frac{\text{No Ans Samples}}{\text{Total Samples}}$ Wrong AnsTotal $\frac{\text{Wrong Ans}}{\text{Total Ans}}$ Correct Ans (NR) Correct $\frac{\text{Correct Ans (NR) }}{\text{Correct Ans}}$ Wrong Ans (NR)Wrong $\frac{\text{Wrong Ans (NR)}}{\text{Wrong Ans}}$ No Ans Samples Total $\frac{\text{No Ans Samples }}{\text{Total Samples}}$ Wrong Ans (NR)Total $\frac{\text{Wrong Ans (NR)}}{\text{Total Ans}}$ Wrong Ans (R)Total $\frac{\text{Wrong Ans (R)}}{\text{Total Ans}}$ WebQSP RoG-Joint 0/1559 = 0% 4605/10206 = 45% 170/5601 = 3% 1178/4605 = 26% 0/69 = 0% 24/126 = 19% 70/126 = 56% RoG-Sep 0/1559 = 0% 6867/12401 = 55% 327/5534 = 6% 2494/6867 = 36% 0/69 = 0% 71/162 = 44% 43/162 = 27% SubgraphRAG (Llama3.1-8B) 29/1559 = 2% 1397/5850 = 24% 59/4453 = 1% 84/1397 = 6% 13/69 = 19% 16/107 = 15% 71/107 = 66% SubgraphRAG (GPT4o-mini) 12/1559 = 1% 1802/8011 = 22% 37/6209 = 1% 126/1802 = 7% 7/69 = 10% 17/83 = 20% 31/83 = 37% CWQ RoG-Joint 0/2848 = 0% 3651/6709 = 54% 305/3058 = 10% 1293/3651 = 35% 0/683 = 0% 1835/2729 = 67% 363/2729 = 13% RoG-Sep 0/2848 = 0% 3183/6129 = 52% 341/2946 = 12% 1295/3183 = 41% 0/683 = 0% 1840/2620 = 70% 268/2620 = 10% SubgraphRAG (Llama3.1-8B) 210/2848 = 7% 2417/5028 = 48% 21/2611 = 1% 98/2417 = 4% 199/683 = 29% 114/1052 = 11% 819/1052 = 78% SubgraphRAG (GPT4o-mini) 203/2848 = 7% 2194/5205 = 42% 30/3011 = 1% 159/2194 = 7% 137/683 = 20% 170/953 = 18% 564/953 = 59%